12kg of a fluid per minutes goes through a reversible steady flow process. The properties of the fluid at inlet are p1=1.4bar, P1=25kg\m3 C1= 120m\s, U1=920kj\kg and at the exit p2=5.6kg\m3,C2=180m\s,P25kg\m3,U2=720kg\kg. During the passage the fluid rejects 60m. Determine the change in enthalpy and work done during the process
1
Expert's answer
2020-06-29T14:04:50-0400

Solution.

"m=12kg\/min=0.2kg\/s;"

"P_1=1bar=10^5Pa;"

"\rho_1=25kg/m^3;"

"C_1=120m\/s;"

"u_1=920kJ/kg=920\sdot10^3J/kg;"

"P_2=5.6bar=5.6\sdot10^5Pa;;"

"\rho_2=5kg/m^3;"

"C_2=180m\/s;"

"u_2=720kJ/kg=720\sdot10^3J/kg;"

"Q=60kJ/s=60\sdot10^3J/s;"

"\Delta Z=60m;"

"\Delta h=\Delta u+\Delta(pv);"

"\Delta (pv)=\dfrac{P_2v_2-P_1v_1}{1}=\dfrac{P_2}{\rho_2}-\dfrac{P_1}{\rho_1};"

"\Delta h=(720-920)\sdot10^3+\dfrac{5.6\sdot10^5}{5}-\dfrac{10^5}{25}=-88.48\sdot10^3J/kg=-88.48kJ/kg;"

"Q=\Delta KE+\Delta PE+\Delta h+W;"

"\Delta KE=\dfrac{C_2^2-C_1^2}{2};"

"\Delta KE=\dfrac{180^2-120^2}{2}=9000J=9kJ" or "=9kJ\/kg;"

"\Delta PE=\Delta Zg;"

"\Delta PE=60\sdot9.81=588.6J =0.588kJ" or "=0.588kJ\/kg;"

"Q=\dfrac{60\sdot10^3}{0.2}=300kJ/kg;"

"-300=9+0.588-88.48+W;"

"W=-221.1kJ\/kg=-221.1kW;"

Answer: "\Delta h=-88.48kJ/kg;"

"W=-221.1kW."